OpenStudy (anonymous):

The absorbance of 2.52x10^-4 M FeSCN 2+ was found to be .545. The absorbance of a second solution in the same container at the same wavelength and temperature was .318. What was the FeSCN2+ concentration in the second solution?

OpenStudy (anonymous):

I got 4.32x10^-4, but that doesn't work out later in the problem

OpenStudy (anonymous):

The second half of the problem is: The second solution of question 1 was prepared by mixing 10.0 ml of 6.40 x10^-3 M Fe(NO3)3 with 10.0 ml of 4.04x10^-4M KSCN. Calculate the equilibrium concentration of Fe3+ and SCN-, and calculate the equilibrium constant, K, for the reaction: FeSCN 2+ <----> Fe3+ + SCN-

OpenStudy (rogue):

For the concentration of the second solution, I got 1.47 x 10^-4.\[A = \epsilon l c\]\[.545 = 2.52 \times 10^{-4} \times \epsilon l\]\[\epsilon l = 2160\]\[.318 = 2160 \times c\]\[c \approx 1.47 \times 10^{-4} M\]